# A Weibull Mixture Cure Frailty Model for High-dimensional Covariates

### v1

@@ 1 Introduction
When analyzing time-to-event data, the presence of substantial censoring following a prolonged follow-up period often indicates the presence of “long-term survivors” or “cured individuals”, i.e., individuals who may never experience the event of interest. This phenomenon can be observed in certain clinical investigations such as cancer studies, where successful treatment can effectively prevent disease recurrence. This scenario becomes more noticeable among patients who are diagnosed in the early stages of cancer development. The presence of a cure fraction violates a key assumption in traditional survival models, like the Cox Proportional Hazards (PH) model, which assumes that all subjects will inevitably experience the event of interest. When a cure fraction exists, this assumption no longer holds true. Hence, the use of a Cox PH model in such cases can lead to an underestimation of hazard rates and an overestimation of survival probabilities for individuals susceptible to the event [76]. In practice, the cured subjects cannot be observed directly, as they are censored together with the individuals who will eventually experience the outcome. However, the Kaplan-Meier (KM) estimated survival curve can be used for determining the existence of a cure fraction when the follow-up period is sufficiently long. A long and consistent plateau in this curve could imply the existence of a cured subset of subjects. Cure models expand the scope of survival analysis by incorporating a fraction of cured individuals. The mixture cure (MC) model is the most commonly employed cure rate model, which is first introduced by Boag [10] and improved by Berkson and Gage [9]. The MC model incorporates two components: one component represents the cured fraction, i.e., the fraction of individuals who will never experience the event of interest, having a survival probability of one. The other component instead captures the susceptible fraction, i.e., the fraction of individuals for whom the event occurrence is governed by a proper survival distribution, which is often referred to as the “latency distribution”.
@@
In medical and epidemiological studies, it is common to assume the existence of unobserved factors generating heterogeneity among individuals that cannot be explained via the observed covariates. Frailty models can be employed to incorporate and account for unobserved heterogeneity among individuals. These models introduce a random frailty term, often assumed to follow a specific distribution, which captures individual-specific characteristics that affect the hazard or risk of the event. In such way, the frailty term allows a more comprehensive and accurate representation of the survival analysis. Price and Manatunga [76] introduced a gamma frailty term into the latency distribution to address the presence of unobserved risks within the MC model. Peng and Zhang [73] extended this model by incorporating covariates into both the cure rate and the latency distribution, and they named this model the mixture cure gamma frailty model. Due to the presence of missing variables in cure models, Sy and Taylor [84] as well as Peng and Dear [72] employed the EM algorithm [24] to derive maximum likelihood estimates of the model parameters. Moreover, Peng and Zhang [73] studied semi-parametric estimation methods for the mixture cure gamma frailty model using the EM algorithm and a multiple imputation method with low-dimensional covariates. They also examined the identifiability of the general mixture cure frailty model (MCFM) in [74]. Cai et al. [16] developed an R package, called smcure [17], to facilitate the application of semi-parametric estimation techniques to both the proportional hazards MC model and the accelerated failure time MC model.
@@
Thanks to advancements in biomedicine, we can now generate and gather molecular data from diverse modalities, such as genomics, epigenomics, transcriptomics, proteomics, and metabolomics, often resulting in high-dimensional data sets. Due to the volume and complexity of such data, and related challenges and opportunities, more sophisticated methods than the traditional ones are required, especially to deal with variable selection. To this purpose, the use of advanced regularization methods such as least absolute shrinkage and selection operator (lasso) [88], elastic net [111], the adaptive elastic net [112], and others has become a standard practice for enhancing accuracy and interpretation. The three methods mentioned above are established popular choices for performing simultaneous prediction and variable selection in high-dimensional linear regression problems. These techniques allow discovering meaningful patterns, such as identifying relevant prognostic biomarkers for a certain disease, from high-dimensional molecular data. When it comes to generalized linear models, Friedman et al. [33] introduced a novel approach utilizing coordinate descent to compute regularization paths, and also implemented their method in the now very popular R package glmnet [34]. Simon et al. [80] extended the previous work to Cox models for right-censored data, incorporating elastic net regularization, with a very efficient algorithm that was then integrated into the glmnet R package, enabling the solution of significantly larger problems than previously possible. Tay et al. [87] expanded the scope of the elastic net-regularized regression to encompass all generalized linear model families, including Cox models with (start, stop] data and strata, also implemented within the glmnet R package. Adaptive versions of both the lasso [111] and elastic net [112] have also been developed, and their properties explored under specific conditions. In a discussion article, Bühlmann and Meier [13] emphasized the relevance of a multi-step procedure for the adaptive lasso, showing that it allows for sparser models at each step. Furthermore, Xiao and Xu [95] introduced the multi-step adaptive elastic-net, showing via extensive numerical exploration that this method effectively reduces the false positives in the variable selection process while preserving estimation accuracy, and implemented it into the msaenet R package [94]. For further insights into regularization methods and their applications, interested readers may refer to Bühlmann and van de Geer [14], Hastie et al. [39] and James et al. [44].
@@
In the context of survival data within high-dimensional settings, the traditional Cox PH model has been commonly used. Although there has recently been a growing interest in the application of cure models to high-dimensional data, the existing literature on this subject has remained somewhat limited. Liu et al. [60] introduced a variable selection procedure for the semi-parametric MC model using penalties based on the smoothly clipped absolute deviation (SCAD) and lasso. Fan et al. [31] explored the minimax concave penalty (MCP), ridge, and lasso penalties, to estimate the MC model in high-dimensional data sets. Masud et al. [66] studied variable selection problems for both the mixture and promotion time cure models, employing penalty terms such as lasso and adaptive lasso. Baretta and Heuchenne [7] conducted a study on variable selection using SCAD penalties for the semi-parametric PH cure model, considering time-varying covariates, and implemented the method in the penPHcure [8] R package. Sun et al. [83] developed a variable selection methodology that incorporates lasso, adaptive lasso, and SCAD-type penalties, tailored for the semi-parametric promotion time cure model, and particularly suitable for interval-censored data. Bussy et al. [15] proposed a novel statistical model known as C-mix, specifically designed as a mixture-of-experts model for handling censored survival outcomes, to model the patients’ heterogeneity by detecting distinct subgroups, and allowed high-dimensional covariates by employing an elastic net penalization. They also conducted a comparison of the C-mix model with both the MC and Cox PH models through simulation experiments. Shi et al. [78] developed a novel penalty method for the MC model that includes both the MCP and a new penalty term that promotes sign consistency using the covariate effects. Xie and Yu [96] proposed the use of neural networks to address inference in the MC model, demonstrating their favorable predictive ability in high-dimensional settings. Xu et al. [98] considered the variable selection problem by adopting penalties such as lasso, adaptive lasso, and SCAD for the generalized odds rate MC model, particularly in the context of interval-censored data. Lastly, Fu et al. [35] investigated the Weibull MC model, incorporating generalized monotone incremental forward stagewise (GMIFS) [38] and EM algorithm techniques for inference. They applied the EM algorithm to a penalized Weibull MC model, employing a lasso type penalty. Their findings indicated a superior performance of their penalized MC model as compared to alternative methods in various simulation scenarios. It is noteworthy to emphasize that all aforementioned studies on cure models exhibit significant limitations concerning the size of covariates relative to the sample size that can be handled by the model, in both simulation experiments and real-data applications, except for the studies conducted by Bussy et al. [15] and Fu et al. [35].
@@
This study is aimed at developing a novel version of the MCFM useful in instances where high-dimensional covariates are potentially associated with both the cured rate and uncured subjects in the sample. To this purpose, a new penalized EM algorithm is introduced, incorporating adaptive elastic net penalties specifically tailored for the MCFM. In particular, the model assumes that the distribution of uncured subjects follows a Weibull distribution, and that their survival probability at each time depends on a subset of the high-dimensional set of covariates, which is also a target of inference. To the best of our knowledge, this study represents the first exploration of MCFM, i.e. the first attempt at incorporating frailty into the MCM, within high-dimensional settings. Nonetheless, surpassing the predominant focus of the existing literature on MCFM on low-dimensional settings, our study explores scenarios in which the sizes of covariates significantly exceed the sample sizes in both simulated and real-data applications. The rest of this article is organized as follows: in Section 2, we present the classical MCFM and its extension to high-dimensional settings. In Section 3, we introduce the novel penalized MCFM (penMCFM), with the associated adaptive EM algorithm for inference, in details. The performances of our proposed methods are compared with some competitors through a comprehensive Monte Carlo simulation study in Section 4. In Section 5, all considered methods are applied to publicly available breast cancer RNA-seq data, with the aim of identifying potential biomarker genes. We also conduct functional enrichment analyses for better interpretation, and determine a prognostic risk score for enhanced validation, of all the identified biomarker genes. Finally, in Section 6 we provide a brief summary and discussion of the study, also mentioning potential future research directions.
@@ 2.1 Classical Mixture Cure Frailty Models
where $\pi(\mathbf{z})$ is the probability of a subject being uncured conditionally on $\mathbf{z}$, and $S(t|Y=1,\mathbf{x})$ is the survival function of the lifetime distribution of uncured subjects conditionally on $\mathbf{x}$. When the mixture cure model defined in (2) is specified via proportional hazards, we get the following PH mixture cure model
$S_{pop}(t|\mathbf{x,z})=1-\pi(\mathbf{z})+\pi(\mathbf{z})S_{0}(t)\exp(\bm{\beta}\mathbf{x}),$ (3)
where $S_{0}(t)$ is the baseline survival function, and $\bm{\beta}$ the vector of unknown regression coefficients, for the latency distribution.
In medical and epidemiological studies, frailty models can be considered as an extension of PH models bringing the modeling closer to the data complexity. A frailty is defined as an unobservable, random, multiplicative factor acting on the hazard function. Let $W_{i}$ be a non-negative frailty random variable associated to the $i$th subject, with cumulative distribution function (cdf) $F_{W_{i}}(w)$. The hazard function of the $i$th subject with frailty $W_{i}$ is
$h(t|W_{i})=W_{i}h_{0}(t)\exp(\bm{\beta}\mathbf{x_{i}}),$
where $h_{0}(t)$ is a baseline hazard function common for all subjects and $\mathbf{x_{i}}$ is a covariate vector for the $i$th subject. If we include the frailty in the latency distribution in model (3), the conditional survival function given the frailty $W$ takes the form
$S(t|Y=1,W,\mathbf{x})=\exp(-We^{\bm{\beta}\mathbf{x}}H_{0}(t)),$
@@
$S_{pop}(t|\mathbf{x,z})=1-\pi(\mathbf{z})+\pi(\mathbf{z})L_{W}(e^{\bm{\beta}\mathbf{x}}H_{0}(t)),$ (4)
Finally, concerning the modeling of the effect of the covariates $\mathbf{z}$ on the incidence, as previously proposed in [32] we use a logistic regression model of the form $\pi(\mathbf{z})=e^{\mathbf{bz}}/(1+e^{\mathbf{bz}}),$ where $\mathbf{b}$ is a vector of unknown regression coefficients. Model (4) was first introduced by Peng and Zhang [73] and called the mixture cure frailty model. Model (4) reduces to the PH mixture cure model in (3) when there is no frailty effect, namely $W\equiv 1$, and it reduces to a standard frailty model when there is no cure fraction existing in the population, namely $\pi(\mathbf{z})\equiv 1$.
@@
$S_{pop}(t|\mathbf{x,z})$ $=$ $\frac{1}{1+e^{\mathbf{bz}}}+\frac{e^{\mathbf{bz}}}{1+e^{\mathbf{bz}}}\left(1+\frac{\alpha t^{\gamma}e^{\bm{\beta}\mathbf{x}}}{\theta}\right)^{-\theta}.$ (5)
@@ 2.2 Mixture Cure Frailty Models with high-dimensional covariates
In the context of medical applications, and particularly in cancer studies, molecular data of several kinds can be used to better estimate both the cured fraction, and the uncured patients survival. This is particularly relevant in cancer studies, since tissue samples originating from biopsies of the tumor are often available, and these can provide excellent information of the tumor composition and characterization. Such molecular omics data often include genomics, epigenomics, transcriptomics, proteomics, metabolomics and radiomics information, often generated using advanced high-throughput technologies to allow the understanding of complex biological systems and their underlying mechanisms. Common characteristic of all these data layers is their high-dimensionality, meaning that the number of variables $p$ included in each omics data layer is much larger than the sample size $n$. We plan to modify the mixture cure frailty model as discussed above so that high-dimensional omics data can be used as covariates in the model.

### v2

@@ 1 Introduction
When analyzing time-to-event data, the presence of substantial censoring following a prolonged follow-up period often indicates the presence of “long-term survivors” or “cured individuals”, i.e., individuals who may never experience the event of interest. This phenomenon can be observed in certain clinical investigations such as cancer studies, where successful treatment can effectively prevent disease recurrence. This scenario becomes more noticeable among patients who are diagnosed in the early stages of cancer development. The presence of a cured fraction violates a key assumption in traditional survival models, like the Cox Proportional Hazards (PH) model, which assumes that all subjects will inevitably experience the event of interest. When a cured fraction exists, this assumption no longer holds true. Hence, the use of a Cox PH model in such cases can lead to an underestimation of hazard rates and an overestimation of survival probabilities for individuals susceptible to the event [1]. In practice, the cured subjects cannot be observed directly, as they are censored together with the individuals who will eventually experience the outcome. However, the Kaplan-Meier (KM) estimated survival curve can be used for determining the existence of a cured fraction when the follow-up period is sufficiently long. A long and consistent plateau in this curve could imply the existence of a cured subset of subjects. Cure models expand the scope of survival analysis by incorporating a fraction of cured individuals. The mixture cure (MC) model is the most commonly employed cure rate model, which was first introduced by Boag [2] and improved by Berkson and Gage [3]. The MC model incorporates two components: one component represents the cured fraction, i.e., the fraction of individuals who will never experience the event of interest, having a survival probability of one. The other component instead captures the susceptible fraction, i.e., the fraction of individuals for whom the event occurrence is governed by a proper survival distribution, which is often referred to as the “latency distribution”.
@@
In medical and epidemiological studies, it is common to assume the existence of unobserved factors generating heterogeneity among individuals that cannot be explained via the observed covariates. Frailty models can be employed to incorporate and account for unobserved heterogeneity among individuals, to allow for a more accurate modeling of the survival outcome. These models introduce a random frailty term, often assumed to follow a specific distribution, which captures individual-specific characteristics that affect the hazard or risk of the event. Price and Manatunga [1] introduced a gamma frailty term into the latency distribution to address the presence of unobserved risks within the MC model. Peng and Zhang [4] extended this model by incorporating covariates into both the cure rate and the latency distribution, and they named this model the mixture cure gamma frailty model. Due to the presence of missing variables in cure models, Sy and Taylor [5] as well as Peng and Dear [6] employed the EM algorithm [7] to derive maximum likelihood estimates of the model parameters. Moreover, Peng and Zhang [4] studied semi-parametric estimation methods for the mixture cure gamma frailty model using the EM algorithm and a multiple imputation method with low-dimensional covariates. They also examined the identifiability of the general mixture cure frailty model (MCFM) in [8]. Cai et al. [9] developed an R package, called smcure [10], to facilitate the application of semi-parametric estimation techniques to both the proportional hazards MC model and the accelerated failure time MC model. Finally, frailty models have also been used in this context when dealing with correlated data, and particularly to be able to take into account heterogeneity in the presence of recurrent events. For instance, Rondeau et al. [11] studied the analysis of recurrent time-to-event data using frailty models with a cured fraction.
@@
The focus of the present paper is on situations when the available covariates are way more than the subjects included in the sample. Thanks to advancements in biomedicine, we can now generate and gather molecular data from diverse modalities, such as genomics, epigenomics, transcriptomics, proteomics, and metabolomics, often resulting in high-dimensional data sets. Due to the volume and complexity of such data, and related challenges and opportunities, more sophisticated methods than the traditional ones are required, especially to deal with variable selection. To this purpose, the use of advanced regularization methods such as least absolute shrinkage and selection operator (lasso) [12], elastic net [13], the adaptive elastic net [14], and others has become a standard practice for enhancing accuracy and interpretation. The three methods mentioned above are established popular choices for performing simultaneous prediction and variable selection in high-dimensional linear regression problems. These techniques allow discovering meaningful patterns, such as identifying relevant prognostic biomarkers for a certain disease, from high-dimensional molecular data. When it comes to generalized linear models, Friedman et al. [15] introduced a novel approach utilizing coordinate descent to compute regularization paths, and also implemented their method in the now very popular R package glmnet [16]. Simon et al. [17] extended the previous work to Cox models for right-censored data, incorporating elastic net regularization, with a very efficient algorithm that was then integrated into the glmnet R package, enabling the solution of significantly larger problems than previously possible. Tay et al. [18] expanded the scope of the elastic net-regularized regression to encompass all generalized linear model families, including Cox models with (start, stop] data and strata, also implemented within the glmnet R package. Adaptive versions of both the lasso [13] and elastic net [14] have also been developed, and their properties explored under specific conditions. In a discussion article, Bühlmann and Meier [19] emphasized the relevance of a multi-step procedure for the adaptive lasso, showing that it allows for sparser models at each step. Furthermore, Xiao and Xu [20] introduced the multi-step adaptive elastic-net, showing via extensive numerical exploration that this method effectively reduces the false positives in the variable selection process while preserving estimation accuracy, and implemented it into the msaenet R package [21]. For further insights into regularization methods and their applications, interested readers may refer to Bühlmann and van de Geer [22], Hastie et al. [23] and James et al. [24].
@@
In the context of survival data within high-dimensional settings, the traditional Cox PH model has been commonly used. Although there has recently been a growing interest in the application of cure models to high-dimensional data, the existing literature on this subject has remained somewhat limited. Liu et al. [25] introduced a variable selection procedure for the semi-parametric MC model using penalties based on the smoothly clipped absolute deviation (SCAD) and lasso. Fan et al. [26] explored the minimax concave penalty (MCP), ridge, and lasso penalties, to estimate the MC model in high-dimensional data sets. Masud et al. [27] studied variable selection problems for both the mixture and promotion time cure models, employing penalty terms such as lasso and adaptive lasso. Baretta and Heuchenne [28] conducted a study on variable selection using SCAD penalties for the semi-parametric PH cure model, considering time-varying covariates, and implemented the method in the penPHcure [29] R package. Sun et al. [30] developed a variable selection methodology that incorporates lasso, adaptive lasso, and SCAD-type penalties, tailored for the semi-parametric promotion time cure model, and particularly suitable for interval-censored data. Bussy et al. [31] proposed a novel statistical model known as C-mix, specifically designed as a mixture-of-experts model for handling censored survival outcomes, to model the patients’ heterogeneity by detecting distinct subgroups, and allowed high-dimensional covariates by employing an elastic net penalization. They also conducted a comparison of the C-mix model with both the MC and Cox PH models through simulation experiments. Shi et al. [32] developed a novel penalty method for the MC model that includes both the MCP and a new penalty term that promotes sign consistency using the covariate effects. Xie and Yu [33] proposed the use of neural networks to address inference in the MC model, demonstrating their favorable predictive ability in high-dimensional settings. Xu et al. [34] considered the variable selection problem by adopting penalties such as lasso, adaptive lasso, and SCAD for the generalized odds rate MC model, particularly in the context of interval-censored data. Lastly, Fu et al. [35] investigated the Weibull MC model, incorporating generalized monotone incremental forward stagewise (GMIFS) [36] and EM algorithm techniques for inference. They applied the EM algorithm to a penalized Weibull MC model, employing a lasso type penalty. Their findings indicated a superior performance of their penalized MC model as compared to alternative methods in various simulation scenarios. It is noteworthy to emphasize that all aforementioned studies on cure models exhibit significant limitations concerning the size of covariates relative to the sample size that can be handled by the model, in both simulation experiments and real-data applications, with the only exception being the case studies conducted by Bussy et al. [31] and Fu et al. [35]. Moreover, neither of the two latter studies include a frailty term in the model.
@@
This study is therefore aimed at developing a first novel version of the MCFM useful in instances where high-dimensional covariates are potentially associated with both the cured rate and uncured subjects in the sample. To this purpose, a new penalized EM algorithm is introduced, incorporating adaptive elastic net penalties specifically tailored for the MCFM. In particular, the model assumes that the distribution of uncured subjects follows a Weibull distribution, and that their survival probability at each time depends on a subset of the high-dimensional set of covariates, which is also a target of inference. To the best of our knowledge, this study represents the first exploration of MCFM, i.e. the first attempt at incorporating frailty into the MCM, within high-dimensional settings. Nonetheless, surpassing the predominant focus of the existing literature on MCFM on low-dimensional settings, our study explores scenarios in which the sizes of covariates significantly exceed the sample sizes in both simulated and real-data applications. The rest of this article is organized as follows: in Section 2, we present the classical MCFM and its extension to high-dimensional settings. In Section 3, we introduce the novel penalized MCFM (penMCFM), with the associated adaptive EM algorithm for inference, in details. The performances of our proposed methods are compared with some competitors through a comprehensive Monte Carlo simulation study in Section 4. In Section 5, all considered methods are applied to publicly available breast cancer RNA-seq data, with the aim of identifying potential biomarker genes. We also conduct functional enrichment analyses for better interpretation, and determine a prognostic risk score for enhanced validation, of all the identified biomarker genes. Finally, in Section 6 we provide a brief summary and discussion of the study, also mentioning potential future research directions.
@@ 2.1 Classical Mixture Cure Frailty Models
where $\pi(\mathbf{z})$ is the probability of a subject being uncured conditionally on $\mathbf{z}$, and $S(t|Y=1,\mathbf{x})$ is the survival function of the lifetime distribution of uncured subjects conditionally on $\mathbf{x}$. Concerning the modeling of the effect of the covariates $\mathbf{z}$ on the incidence, as previously proposed in [37] we use a logistic regression model of the form $\pi(\mathbf{z})=e^{\mathbf{z}^{\top}\mathbf{b}}/(1+e^{\mathbf{z}^{\top}\mathbf{b}}),$ where $\mathbf{z}^{\top}\in\mathbb{R}^{n\times P_{1}+1}$ is a covariate matrix, with columns $\mathbf{z}_{1},\cdots,\mathbf{z}_{n}\in\mathbb{R}^{P_{1}+1}$, and $\mathbf{b}=(b_{0},b_{1},\cdots,b_{P_{1}})^{\top}\in\mathbb{R}^{P_{1}+1}$ is a vector of unknown regression coefficients. When the mixture cure model defined in (2) is specified via proportional hazards, we get the following PH mixture cure model
$S_{pop}(t|\mathbf{x,z})=1-\pi(\mathbf{z})+\pi(\mathbf{z})S_{0}(t)^{\exp(\mathbf{x}^{\top}\bm{\beta})},$ (3)
where $S_{0}(t)$ is the baseline survival function, $\mathbf{x}^{\top}\in\mathbb{R}^{n\times P_{2}}$ is the covariate matrix, with columns $\mathbf{x_{1}},\cdots,\mathbf{x_{n}}\in\mathbb{R}^{P_{2}}$, and $\bm{\beta}=(\beta_{1},\cdots,\beta_{P_{2}})^{\top}\in\mathbb{R}^{P_{2}}$ is the vector of unknown regression coefficients for the latency distribution.
In medical and epidemiological studies, frailty models extend the Cox PH model to account for unobservable heterogeneity among individuals. This extension provides a more flexible structure for analysis. A frailty is defined as an unobservable, random, multiplicative factor acting on the hazard function. Let $W_{i}$ be a non-negative frailty random variable associated to the $i$th subject, $i=1,\ldots,n$, with cumulative distribution function (cdf) $F_{W_{i}}(w)$. The hazard function of the $i$th subject with frailty $W_{i}$ is
$h(t|W_{i})=W_{i}\;h_{0}(t)\exp(\mathbf{x_{i}}^{\top}\bm{\beta}),$
where $h_{0}(t)$ is a baseline hazard function common for all subjects and $\mathbf{x_{i}}$ is the covariate vector for the $i$th subject. If we include the frailty in the latency distribution in model (3), the conditional survival function given the frailty $W$ takes the form
$S(t|Y=1,W,\mathbf{x})=\exp(-We^{\mathbf{x}^{\top}\bm{\beta}}H_{0}(t)),$
@@
$S_{pop}(t|\mathbf{x,z})=1-\pi(\mathbf{z})+\pi(\mathbf{z})L_{W}(e^{\mathbf{x}^{\top}\bm{\beta}}H_{0}(t)).$ (4)
Model (4) was first introduced by Peng and Zhang [4] and called the mixture cure frailty model. Model (4) reduces to the PH mixture cure model in (3) when there is no frailty effect, namely $W\equiv 1$, and it reduces to a standard frailty model when there is no cure fraction existing in the population, namely $\pi(\mathbf{z})\equiv 1$.
@@
$S_{pop}(t|\mathbf{x,z})$ $=$ $\frac{1}{1+e^{\mathbf{z}^{\top}\mathbf{b}}}+\frac{e^{\mathbf{z}^{\top}\mathbf{b}}}{1+e^{\mathbf{z}^{\top}\mathbf{b}}}\left(1+\frac{\alpha t^{\gamma}e^{\mathbf{x}^{\top}\bm{\beta}}}{\theta}\right)^{-\theta}.$ (5)
@@ 2.2 Mixture Cure Frailty Models with high-dimensional covariates
In the context of medical applications, and particularly in cancer studies, molecular data of several kinds can be used to better estimate both the cured fraction, and the uncured patients survival. This is particularly relevant in cancer studies, since tissue samples originating from biopsies of the tumor are often available, and these can provide excellent information on the tumor composition and characterization. Such molecular omics data often include genomics, epigenomics, transcriptomics, proteomics, metabolomics and radiomics information, often generated using advanced high-throughput technologies to allow the understanding of complex biological systems and their underlying mechanisms. Common characteristic of all these data layers is their high-dimensionality, meaning that the number of variables $p$ included in each omics data layer is much larger than the sample size $n$. We plan to modify the mixture cure frailty model as discussed above so that high-dimensional omics data can be used as covariates in the model.